### v1

@@ Abstract
Domain generalization (DG) aims at solving distribution shift problems in various scenes. Existing approaches are based on Convolution Neural Networks (CNNs) or Vision Transformers (ViTs), which suffer from limited receptive fields or quadratic complexities issues. Mamba, as an emerging state space model (SSM), possesses superior linear complexity and global receptive fields. Despite this, it can hardly be applied to DG to address distribution shifts, due to the hidden state issues and inappropriate scan mechanisms. In this paper, we propose a novel framework for DG, named DGMamba, that excels in strong generalizability toward unseen domains and meanwhile has the advantages of global receptive fields, and efficient linear complexity. Our DGMamba compromises two core components: Hidden State Suppressing (HSS) and Semantic-aware Patch refining (SPR). In particular, HSS is introduced to mitigate the influence of hidden states associated with domain-specific features during output prediction. SPR strives to encourage the model to concentrate more on objects rather than context, consisting of two designs: Prior-Free Scanning (PFS), and Domain Context Interchange (DCI). Concretely, PFS aims to shuffle the non-semantic patches within images, creating more flexible and effective sequences from images, and DCI is designed to regularize Mamba with the combination of mismatched non-semantic and semantic information by fusing patches among domains. Extensive experiments on four commonly used DG benchmarks demonstrate that the proposed DGMamba achieves remarkably superior results to state-of-the-art models. The code will be made publicly available.
@@ 1 Introduction
Motivated by the above facts, we propose DGMamba, a novel State Space Model-based framework for domain generalization that excels in strong generalizability toward unseen domains and meanwhile has the advantages of global receptive fields, and efficient linear complexity. DGMamba comprises two key modules, Hidden State Suppressing (HSS) and Semantic-aware Patch Refining (SPR). Firstly, HSS is presented to eliminate the detrimental effect of non-semantic information contained in hidden states by selectively suppressing the corresponding hidden states during output prediction. By reducing non-semantic information in SSM layers, DGMamba learns domain invariant features. Secondly, SPR is introduced to encourage the model to pay more attention to objects rather than context, consisting of two key designs: Prior-Free Scanning (PFS), and Domain Context Interchange (DCI). Specifically, PFS is designed to shuffle the context patches within images and contribute less to the label prediction. It provides an effective 2D scanning mechanism to traverse 2D images into 1D sequence data. As a result, PFS possesses the ability to shift the model’s attention from the context to the object. Besides, to alleviate the influence of diverse context information and local texture details across different domains, DCI replaces the context patches of images with those from different domains. The proposed DCI brings in local texture noise and regularizes the model on the combination of mismatched context and object. By leveraging both linear complexity and heterogeneous context tokens, DCI learns more robust representation efficiently. Aggregating all these contributions into one architecture, our proposed DGMamba achieves strong results on four DG benchmarks. As shown in Figure 1, compared with previous CNN-based and ViT-based methods, our DGMamba achieves the best trade-off between accuracy and parameters. In summary, we make the following contributions:
@@ 3.2 Hidden States Suppressing
$\bar{C}$ $=C\times M,$
$M$ $=(\Delta A>\alpha)+(1-(\Delta A>\alpha))\Delta A,$
@@ 4 Experiments
Dataset. Following standard protocol in DG [20, 7], we evaluate the effectiveness of our proposed DGMamba and compare it with state-of-the-art methods in DG on four commonly used benchmarks: (1) PACS [36] includes 9991 images categorized into 7 classes exhibiting 4 styles. (2) VLCS [14] involves four datasets, totaling 10729 images distributed in 5 categories. (3) OfficeHome [62] comprises 15588 images in 65 classes from 4 datasets. (4) TerraIncognita [4] encompasses 24330 photographs of 10 kinds of animals taken at 4 diverse locations.
@@
Method Backbone Params. Target domain Avg.($\uparrow$)
@@
GroupDRO [54] (ICLR’2019) ResNet50 23M 83.5 79.1 96.7 78.3 84.4
VREx [34] (ICML’2021) ResNet50 23M 86.0 79.1 96.9 77.7 84.9
RSC [29] (ECCV’2020) ResNet50 23M 85.4 79.7 97.6 78.2 85.2
MTL [6] (JMLR’2021) ResNet50 23M 87.5 77.1 96.4 77.3 84.6
Mixstyle [90] (ICLR’2021) ResNet50 23M 86.8 79.0 96.6 78.5 85.2
SagNet [49] (CVPR’2021) ResNet50 23M 87.4 80.7 97.1 80.0 86.3
ARM [78] (NeurIPS’2021) ResNet50 23M 86.8 76.8 97.4 79.3 85.1
SWAD [7] (NeurIPS’2021) ResNet50 23M 89.3 83.4 97.3 82.5 88.1
PCL [73] (CVPR’2022) ResNet50 23M 90.2 83.9 98.1 82.6 88.7
SAGM [64] (CVPR’2023) ResNet50 23M 87.4 80.2 98.0 80.8 86.6
iDAG [30] (ICCV’2023) ResNet50 23M 90.8 83.7 98.0 82.7 88.8
GMDG [60] (CVPR’2024) ResNet50 23M 84.7 81.7 97.5 80.5 85.6
SDViT [59] (ACCV’2022) DeiT-S 22M 87.6 82.4 98.0 77.2 86.3
GMoE [35] (ICLR’2023) DeiT-S 34M 89.4 83.9 99.1 74.5 86.7
DGMamba (ours) VMamba-T 22M 91.3 87.0 99.0 87.3 91.2
@@
Table 2: Results on VLCS with our DGMamba. The best results are bolded.
Method Backbone Params. Target domain Avg.($\uparrow$)
Caltech LabelMe SUN PASCAL
GroupDRO [54](ICLR’2019) ResNet50 23M 97.3 63.4 69.5 76.7 76.7
VREx [34] (ICML’2021) ResNet50 23M 98.4 64.4 74.1 76.2 78.3
RSC [29] (ECCV’2020) ResNet50 23M 97.9 62.5 72.3 75.6 77.1
MTL [6] (JMLR’2021) ResNet50 23M 97.8 64.3 71.5 75.3 77.2
Mixstyle [90] (ICLR’2021) ResNet50 23M 98.6 64.5 72.6 75.7 77.9
SagNet [49] (CVPR’2021) ResNet50 23M 97.9 64.5 71.4 77.5 77.8
ARM [78] (NeurIPS’2021) ResNet50 23M 98.7 63.6 71.3 76.7 77.6
SWAD [7] (NeurIPS’2021) ResNet50 23M 98.8 63.3 75.3 79.2 79.1
PCL [73] (CVPR’2022) ResNet50 23M 99.0 63.6 73.8 75.6 78.0
SAGM [64] (CVPR’2023) ResNet50 23M 99.0 65.2 75.1 80.7 80.0
iDAG [30] (ICCV’2023) ResNet50 23M 98.1 62.7 69.9 77.1 76.9
GMDG [60] (CVPR’2024) ResNet50 23M 98.3 65.9 73.4 79.3 79.2
SDViT [59] (ACCV’2022) DeiT-S 22M 96.8 64.2 76.2 78.5 78.9
GMoE [35] (ICLR’2023) DeiT-S 34M 96.9 63.2 72.3 79.5 78.0
DGMamba (ours) VMamba-T 22M 98.9 64.3 79.2 80.8 80.8
@@ 4.1 Main results
Method Backbone Params. Target domain Avg.($\uparrow$)
Art Clipart Product Real
GroupDRO [54] (ICLR’2019) ResNet50 23M 60.4 52.7 75.0 76.0 66.0
VREx [34] (ICML’2021) ResNet50 23M 60.7 53.0 75.3 76.6 66.4
RSC [29] (ECCV’2020) ResNet50 23M 60.7 51.4 74.8 75.1 65.5
MTL [6] (JMLR’2021) ResNet50 23M 61.5 52.4 74.9 76.8 66.4
Mixstyle [90] (ICLR’2021) ResNet50 23M 51.1 53.2 68.2 69.2 60.4
SagNet [49] (CVPR’2021) ResNet50 23M 63.4 54.8 75.8 78.3 68.1
ARM [78] (NeurIPS’2021) ResNet50 23M 58.9 51.0 74.1 75.2 64.8
SWAD [7] (NeurIPS’2021) ResNet50 23M 66.1 57.7 78.4 80.2 70.6
PCL [73] (CVPR’2022) ResNet50 23M 67.3 59.9 78.7 80.7 71.6
SAGM [64] (CVPR’2023) ResNet50 23M 65.4 57.0 78.0 80.0 70.1
iDAG [30] (ICCV’2023) ResNet50 23M 68.2 57.9 79.7 81.4 71.8
GMDG [60] (CVPR’2024) ResNet50 23M 68.9 56.2 79.9 82.0 70.7
SDViT [59] (ACCV’2022) DeiT-S 22M 68.3 56.3 79.5 81.8 71.5
GMoE [35] (ICLR’2023) DeiT-S 34M 69.3 58.0 79.8 82.6 72.4
DGMamba (ours) VMamba-T 22M 76.2 61.8 83.9 86.1 77.0
Method Backbone Params. Target domain Avg.($\uparrow$)
L100 L38 L43 L46
GroupDRO [54] (ICLR’2019) ResNet50 23M 41.2 38.6 56.7 36.4 43.2
VREx [34] (ICML’2021) ResNet50 23M 48.2 41.7 56.8 38.7 46.4
RSC [29] (ECCV’2020) ResNet50 23M 50.2 39.2 56.3 40.8 46.6
MTL [6] (JMLR’2021) ResNet50 23M 49.3 39.6 55.6 37.8 45.6
Mixstyle [90] (ICLR’2021) ResNet50 23M 54.3 34.1 55.9 31.7 44.0
SagNet [49] (CVPR’2021) ResNet50 23M 53.0 43.0 57.9 40.4 48.6
ARM [78] (NeurIPS’2021) ResNet50 23M 49.3 38.3 55.8 38.7 45.5
SWAD [7] (NeurIPS’2021) ResNet50 23M 55.4 44.9 59.7 39.9 50.0
PCL [73] (CVPR’2022) ResNet50 23M 58.7 46.3 60.0 43.6 52.1
SAGM [64] (CVPR’2023) ResNet50 23M 54.8 41.4 57.7 41.3 48.8
iDAG [30] (ICCV’2023) ResNet50 23M 58.7 35.1 57.5 33.0 46.1
GMDG [60] (CVPR’2024) ResNet50 23M 59.8 45.3 57.1 38.2 50.1
SDViT [59] (ACCV’2022) DeiT-S 22M 55.9 31.7 52.2 37.4 44.3
GMoE [35] (ICLR’2023) DeiT-S 34M 59.2 34.0 50.7 38.5 45.6
DGMamba (ours) VMamba-T 22M 62.7 48.3 61.1 46.4 54.6
@@ 4.2 Ablation study and analysis
Table 7: Comparison of the proposed Hidden State Suppressing (HSS) with Hidden State Masking (HSM) on PACS using recognition accuracy (%). Threshold 0 0.15 0.2 0.25 HSM 89.4 90.2 90.3 90.1 HSS 89.4 90.1 90.5 90.8
@@
iDAG [30] (ICCV’2023) ResNet50 23 8 94 88.8
iDAG [30] (ICCV’2023) ResNet101 41 15 495 89.2
@@
VMamba [44] (Arxiv’2024) VMamba-T 22 5 225 89.4
DGMamba (ours) VMamba-T 22 5 233 91.2
@@ 5 Conclusion
This work is the first endeavor to explore the generalizability of the SSM-based model (Mamba) in DG. We propose a novel framework named DGMamba, that contains two pivotal techniques. Firstly, we design a novel Hidden State Suppressing to alleviate the adverse effect of domain-specific information conveyed in hidden states. Secondly, we propose Semantic-aware Patch Refining, which consists of Prior-Free Scanning and Domain Context Interchange. Both aim to direct the model’s focus toward the object rather than the context. Extensive experiments on four widely used DG benchmarks show the superiority of DGMamba compared with SOTA DG methods based on CNN or ViT. We believe our work builds a solid baseline for exploiting SSMs for the DG community.

### v2

@@ Abstract
Domain generalization (DG) aims at solving distribution shift problems in various scenes. Existing approaches are based on Convolution Neural Networks (CNNs) or Vision Transformers (ViTs), which suffer from limited receptive fields or quadratic complexities issues. Mamba, as an emerging state space model (SSM), possesses superior linear complexity and global receptive fields. Despite this, it can hardly be applied to DG to address distribution shifts, due to the hidden state issues and inappropriate scan mechanisms. In this paper, we propose a novel framework for DG, named DGMamba, that excels in strong generalizability toward unseen domains and meanwhile has the advantages of global receptive fields, and efficient linear complexity. Our DGMamba compromises two core components: Hidden State Suppressing (HSS) and Semantic-aware Patch refining (SPR). In particular, HSS is introduced to mitigate the influence of hidden states associated with domain-specific features during output prediction. SPR strives to encourage the model to concentrate more on objects rather than context, consisting of two designs: Prior-Free Scanning (PFS), and Domain Context Interchange (DCI). Concretely, PFS aims to shuffle the non-semantic patches within images, creating more flexible and effective sequences from images, and DCI is designed to regularize Mamba with the combination of mismatched non-semantic and semantic information by fusing patches among domains. Extensive experiments on five commonly used DG benchmarks demonstrate that the proposed DGMamba achieves remarkably superior results to state-of-the-art models. The code will be made publicly available.
@@ 1 Introduction
Motivated by the above facts, we propose DGMamba, a novel State Space Model-based framework for domain generalization that excels in strong generalizability toward unseen domains and meanwhile has the advantages of global receptive fields, and efficient linear complexity. DGMamba comprises two key modules, Hidden State Suppressing (HSS) and Semantic-aware Patch Refining (SPR). Firstly, HSS is presented to eliminate the detrimental effect of non-semantic information contained in hidden states by selectively suppressing the corresponding hidden states during output prediction. By reducing non-semantic information in SSM layers, DGMamba learns domain invariant features. Secondly, SPR is introduced to encourage the model to pay more attention to objects rather than context, consisting of two key designs: Prior-Free Scanning (PFS), and Domain Context Interchange (DCI). Specifically, PFS is designed to shuffle the context patches within images and contribute less to the label prediction. It provides an effective 2D scanning mechanism to traverse 2D images into 1D sequence data. As a result, PFS possesses the ability to shift the model’s attention from the context to the object. Besides, to alleviate the influence of diverse context information and local texture details across different domains, DCI replaces the context patches of images with those from different domains. The proposed DCI brings in local texture noise and regularizes the model on the combination of mismatched context and object. By leveraging both linear complexity and heterogeneous context tokens, DCI learns more robust representation efficiently. Aggregating all these contributions into one architecture, our proposed DGMamba achieves strong results on five DG benchmarks. As shown in Figure 1, compared with previous CNN-based and ViT-based methods, our DGMamba achieves the best trade-off between accuracy and parameters. In summary, we make the following contributions:
@@ 3.2 Hidden States Suppressing
$\bar{C}$ $=C\odot M,$
$M$ $=(\Delta A>\alpha)+(1-(\Delta A>\alpha))\odot\Delta A,$
@@ 4 Experiments
Dataset. Following standard protocol in DG [20, 7], we evaluate the effectiveness of our proposed DGMamba and compare it with state-of-the-art methods in DG on five commonly used benchmarks: (1) PACS [36] includes 9991 images categorized into 7 classes exhibiting 4 styles. (2) VLCS [14] involves four datasets, totaling 10729 images distributed in 5 categories. (3) OfficeHome [63] comprises 15588 images in 65 classes from 4 datasets. (4) TerraIncognita [4] encompasses 24330 photographs of 10 kinds of animals taken at 4 diverse locations. (5) DomainNet [52] comprises 586575 images categorized into 345 classes from six domains.
@@
Method Backbone Venue Target domain Avg.($\uparrow$)
@@
GroupDRO [55] ResNet50 ICLR’2019 83.5 79.1 96.7 78.3 84.4
VREx [34] ResNet50 ICML’2021 86.0 79.1 96.9 77.7 84.9
RSC [29] ResNet50 ECCV’2020 85.4 79.7 97.6 78.2 85.2
MTL [6] ResNet50 JMLR’2021 87.5 77.1 96.4 77.3 84.6
Mixstyle [91] ResNet50 ICLR’2021 86.8 79.0 96.6 78.5 85.2
SagNet [49] ResNet50 CVPR’2021 87.4 80.7 97.1 80.0 86.3
ARM [79] ResNet50 NeurIPS’2021 86.8 76.8 97.4 79.3 85.1
SWAD [7] ResNet50 NeurIPS’2021 89.3 83.4 97.3 82.5 88.1
PCL [74] ResNet50 CVPR’2022 90.2 83.9 98.1 82.6 88.7
SAGM [65] ResNet50 CVPR’2023 87.4 80.2 98.0 80.8 86.6
iDAG [30] ResNet50 ICCV’2023 90.8 83.7 98.0 82.7 88.8
GMDG [61] ResNet50 CVPR’2024 84.7 81.7 97.5 80.5 85.6
SDViT [60] DeiT-S ACCV’2022 87.6 82.4 98.0 77.2 86.3
GMoE [35] DeiT-S ICLR’2023 89.4 83.9 99.1 74.5 86.7
DGMamba VMamba-T - 91.3 87.0 99.0 87.3 91.2
Table 2: Results on VLCS with our DGMamba. The best results are bolded.
Method Backbone Venue Target domain Avg.($\uparrow$)
Caltech LabelMe SUN PASCAL
GroupDRO [55] ResNet50 ICLR’2019 97.3 63.4 69.5 76.7 76.7
VREx [34] ResNet50 ICML’2021 98.4 64.4 74.1 76.2 78.3
RSC [29] ResNet50 ECCV’2020 97.9 62.5 72.3 75.6 77.1
MTL [6] ResNet50 JMLR’2021 97.8 64.3 71.5 75.3 77.2
Mixstyle [91] ResNet50 ICLR’2021 98.6 64.5 72.6 75.7 77.9
SagNet [49] ResNet50 CVPR’2021 97.9 64.5 71.4 77.5 77.8
ARM [79] ResNet50 NeurIPS’2021 98.7 63.6 71.3 76.7 77.6
SWAD [7] ResNet50 NeurIPS’2021 98.8 63.3 75.3 79.2 79.1
PCL [74] ResNet50 CVPR’2022 99.0 63.6 73.8 75.6 78.0
SAGM [65] ResNet50 CVPR’2023 99.0 65.2 75.1 80.7 80.0
iDAG [30] ResNet50 ICCV’2023 98.1 62.7 69.9 77.1 76.9
GMDG [61] ResNet50 CVPR’2024 98.3 65.9 73.4 79.3 79.2
SDViT [60] DeiT-S ACCV’2022 96.8 64.2 76.2 78.5 78.9
GMoE [35] DeiT-S ICLR’2023 96.9 63.2 72.3 79.5 78.0
DGMamba VMamba-T - 98.9 64.3 79.2 80.8 80.8
@@
Method Backbone Venue Target domain Avg.($\uparrow$)
Art Clipart Product Real
GroupDRO [55] ResNet50 ICLR’2019 60.4 52.7 75.0 76.0 66.0
VREx [34] ResNet50 ICML’2021 60.7 53.0 75.3 76.6 66.4
RSC [29] ResNet50 ECCV’2020 60.7 51.4 74.8 75.1 65.5
MTL [6] ResNet50 JMLR’2021 61.5 52.4 74.9 76.8 66.4
Mixstyle [91] ResNet50 ICLR’2021 51.1 53.2 68.2 69.2 60.4
SagNet [49] ResNet50 CVPR’2021 63.4 54.8 75.8 78.3 68.1
ARM [79] ResNet50 NeurIPS’2021 58.9 51.0 74.1 75.2 64.8
SWAD [7] ResNet50 NeurIPS’2021 66.1 57.7 78.4 80.2 70.6
PCL [74] ResNet50 CVPR’2022 67.3 59.9 78.7 80.7 71.6
SAGM [65] ResNet50 CVPR’2023 65.4 57.0 78.0 80.0 70.1
iDAG [30] ResNet50 ICCV’2023 68.2 57.9 79.7 81.4 71.8
GMDG [61] ResNet50 CVPR’2024 68.9 56.2 79.9 82.0 70.7
SDViT [60] DeiT-S ACCV’2022 68.3 56.3 79.5 81.8 71.5
GMoE [35] DeiT-S ICLR’2023 69.3 58.0 79.8 82.6 72.4
DGMamba VMamba-T - 76.2 61.8 83.9 86.1 77.0
Method Backbone Venue Target domain Avg.($\uparrow$)
L100 L38 L43 L46
GroupDRO [55] ResNet50 ICLR’2019 41.2 38.6 56.7 36.4 43.2
VREx [34] ResNet50 ICML’2021 48.2 41.7 56.8 38.7 46.4
RSC [29] ResNet50 ECCV’2020 50.2 39.2 56.3 40.8 46.6
MTL [6] ResNet50 JMLR’2021 49.3 39.6 55.6 37.8 45.6
Mixstyle [91] ResNet50 ICLR’2021 54.3 34.1 55.9 31.7 44.0
SagNet [49] ResNet50 CVPR’2021 53.0 43.0 57.9 40.4 48.6
ARM [79] ResNet50 NeurIPS’2021 49.3 38.3 55.8 38.7 45.5
SWAD [7] ResNet50 NeurIPS’2021 55.4 44.9 59.7 39.9 50.0
PCL [74] ResNet50 CVPR’2022 58.7 46.3 60.0 43.6 52.1
SAGM [65] ResNet50 CVPR’2023 54.8 41.4 57.7 41.3 48.8
iDAG [30] ResNet50 ICCV’2023 58.7 35.1 57.5 33.0 46.1
GMDG [61] ResNet50 CVPR’2024 59.8 45.3 57.1 38.2 50.1
SDViT [60] DeiT-S ACCV’2022 55.9 31.7 52.2 37.4 44.3
GMoE [35] DeiT-S ICLR’2023 59.2 34.0 50.7 38.5 45.6
DGMamba VMamba-T - 62.7 48.3 61.1 46.4 54.6
@@ 4.1 Main results
Results on DomainNet datasets. To further assess the efficacy of mitigating distribution shifts in large-scale benchmarks, we test the proposed DGMamba on DomainNet [52] and compare it with state-of-the-art DG methods. As illustrated by Table 5, our proposed DGMamba demonstrates a substantial improvement of 2.7% compared to the state-of-the-art approach in terms of average generalization performance across diverse domains. Remarkably, our DGMamba attains the best performance in five out of the six domains. These findings underscore the superiority of the proposed DGMamba in tackling distribution shifts in real-world applications.
Table 5: Results on DomainNet with our DGMamba. The best results are bolded.
Method Backbone Venue Target domain Avg.($\uparrow$)
Clipart Infograph painting Quickdraw Real Sketch
GroupDRO [55] ResNet50 ICLR’2019 47.2 17.5 33.8 9.3 51.6 40.1 33.3
VREx [34] ResNet50 ICML’2021 47.3 16.0 35.8 10.9 49.6 42.0 33.6
RSC [29] ResNet50 ECCV’2020 55.0 18.3 44.4 12.2 55.7 47.8 38.9
MTL [6] ResNet50 JMLR’2021 57.9 18.5 46.0 12.5 59.5 49.2 40.6
Mixstyle [91] ResNet50 ICLR’2021 51.9 13.3 37.0 12.3 46.1 43.4 34.0
SagNet [49] ResNet50 CVPR’2021 57.7 19.0 45.3 12.7 58.1 48.8 40.3
ARM [79] ResNet50 NeurIPS’2021 49.7 16.3 40.9 9.4 53.4 43.5 35.5
SWAD [7] ResNet50 NeurIPS’2021 66.0 22.4 53.5 16.1 65.8 55.5 46.5
PCL [74] ResNet50 CVPR’2022 67.9 24.3 55.3 15.7 66.6 56.4 47.7
SAGM [65] ResNet50 CVPR’2023 64.9 21.1 51.5 14.8 64.1 53.6 45.0
iDAG [30] ResNet50 ICCV’2023 67.9 24.2 55.0 16.4 66.1 56.9 47.7
GMDG [61] ResNet50 CVPR’2024 63.4 22.4 51.4 13.4 64.4 52.4 44.6
SDViT [60] DeiT-S ACCV’2022 63.4 22.9 53.7 15.0 67.4 52.6 45.8
GMoE [35] DeiT-S ICLR’2023 68.2 24.7 55.7 16.3 69.1 55.4 48.3
DGMamba VMamba-T - 67.0 27.9 56.5 18.4 69.5 57.9 49.6
@@ 4.2 Ablation study and analysis
Table 8: Comparison of the proposed Hidden State Suppressing (HSS) with Hidden State Masking (HSM) on PACS using recognition accuracy (%).
Threshold 0 0.15 0.2 0.25
HSM 89.4 90.2 90.3 90.1
HSS 89.4 90.1 90.5 90.8
@@
iDAG [30] (ICCV’2023) ResNet50 25 8 94 88.8
iDAG [30] (ICCV’2023) ResNet101 43 15 495 89.2
@@
VMamba [44] (Arxiv’2024) VMamba-T 31 5 225 89.4
DGMamba (ours) VMamba-T 31 5 233 91.2
@@ 5 Conclusion
This work is the first endeavor to explore the generalizability of the SSM-based model (Mamba) in DG. We propose a novel framework named DGMamba, that contains two pivotal techniques. Firstly, we design a novel Hidden State Suppressing to alleviate the adverse effect of domain-specific information conveyed in hidden states. Secondly, we propose Semantic-aware Patch Refining, which consists of Prior-Free Scanning and Domain Context Interchange. Both aim to direct the model’s focus toward the object rather than the context. Extensive experiments on five widely used DG benchmarks show the superiority of DGMamba compared with SOTA DG methods based on CNN or ViT. We believe our work builds a solid baseline for exploiting SSMs for the DG community.
@@
Appendix A Supplementary Material
A.1 Experiment Details
Dataset Details. In this section, we provide details of the 5 commonly used DG benchmarks in Table 10. The data in these datasets originates from various sources with distinct characteristics, such as hand-drawn illustrations, software-composited images, object-centered photographs, and scene-centered shots.
Evaluation Protocols. Following the standard evaluation protocols [20, 92, 7, 81, 46], we report the generalization performance based on the train-domain validation, i.e., selecting one domain as the target domain and training on the remaining domains. Each source domain is split with an 80%/20% train/validation ratio. The validation parts from these source domains collectively constitute the validation set used for the model evaluation.
Hyperparameters for DomainNet. The benchmark DomainNet [52] presents a great challenge, containing 586575 images. Training 10,000 iterations represents less than two complete epochs on DomainNet, which is insufficient for the model to converge effectively. Consequently, recent state-of-the-art DG works [7, 35, 82] increase the iterations to 15000 with a batch size of 32 for each domain. To make a fair comparison while considering the constraints posed by GPU, our proposed DGMamba undergoes 80000 iterations with a batch size of 6 for each source domain. The corresponding initial learning rate is searched in [7e-4, 8.5e-4].
Table 10: Statistics of DG benchmarks.
Dataset Domain # image # image # class
PACS [36] Art 2048 9991 7
Photo 1670
Clipart 2344
Sketch 3929
VLCS [14] Caltech 1415 10729 5
LabelMe 2656
SUN 3282
PASCAL 3376
OfficeHome [63] Art 2427 15588 65
Clipart 4365
Product 4439
Real 4357
TerraIncognita [4] L100 4741 24330 10
L38 9736
L43 3970
L46 5883
DomainNet [52] Clipart 48129 586575 345
Infograph 51605
Painting 72266
Quickdraw 172500
Real 172947
Sketch 69128
Overall Architecture. For an input image $x$, it undergoes initial processing into patches $x_{i}\in x$ utilizing a convolution neural network with layer normalization. These patches are subsequently scanned to be processed by four Mamba blocks. Down-sampling is applied after the first three Mamba blocks to generate feature maps with different resolutions. Finally, prediction is performed using a linear classifier subsequent to the layer normalization, average pooling, and flattening operations.
A.2 Additional Experiments
Comparison across Five Benchmarks. In Table 11, we present a summary of the generalization performance results across five DG benchmarks with training-validation selection. Notably, our proposed DGMamba remarkably outperforms the state-of-the-art DG approaches across all benchmarks, yielding an average gain of 4.4%. The significant enhancements across diverse scenarios from different benchmarks underscore the excellent ability of our DGMamba to tackle distribution shifts.
Table 11: Comparison of state-of-the-art DG methods with our DGMamba. Out-of-domain generalization performance on five commonly used benchmarks is reported. The best results are bolded.
Method Backbone Venue Dataset Avg.($\uparrow$)
PACS VLCS OfficeHome TerraIncognita DomainNet
VREx [34] ResNet50 ICML’2021 84.9 78.3 66.4 46.4 33.6 61.9
RSC [29] ResNet50 ECCV’2020 85.2 77.1 65.5 46.6 38.9 62.7
MTL [6] ResNet50 JMLR’2021 84.6 77.2 66.4 45.6 40.6 62.9
Mixstyle [91] ResNet50 ICLR’2021 85.2 77.9 60.4 44.0 34.0 60.3
SagNet [49] ResNet50 CVPR’2021 86.3 77.8 68.1 48.6 40.3 64.2
ARM [79] ResNet50 NeurIPS’2021 85.1 77.6 64.8 45.5 35.5 61.7
SWAD [7] ResNet50 NeurIPS’2021 88.1 79.1 70.6 50.0 46.5 66.9
PCL [74] ResNet50 CVPR’2022 88.7 78.0 71.6 52.1 47.7 67.6
SAGM [65] ResNet50 CVPR’2023 86.6 80.0 70.1 48.8 45.0 66.1
iDAG [30] ResNet50 ICCV’2023 88.8 76.9 71.8 46.1 47.7 66.3
GMDG [61] ResNet50 CVPR’2024 85.6 79.2 70.7 50.1 44.6 66.0
SDViT [60] DeiT-S ACCV’2022 86.3 78.9 71.5 44.3 45.8 65.4
GMoE [35] DeiT-S ICLR’2023 86.7 78.0 72.4 45.6 48.3 66.2
DGMamba VMamba-T Ours 91.2 80.8 77.0 54.6 49.6 70.6
Figure 7: Visualizations with t-SNE embeddings [62] illustrating features’ distribution gaps between the source and target domains generated by (a) iDAG [30], (b) GMoE [35], (c) VMamba [44], and (d) DGMamba (ours), respectively. Our proposed DGMamba displays the superior feature alignment.
Visualization for Distribution Gaps. To visually demonstrate that our proposed DGMamba could effectively address the distribution shifts between diverse domains, we employ the t-SNE technique [62] to examine the representation’s distribution gaps across domains. We conduct experiments on PACS with ‘Art’ as the target domain. Figure 7 presents the visualization results based on the CNN-based method iDAG [30], ViT-based method GMoE [35], VMamba [44], and our proposed DGMamba, respectively. Notably, our DGMamba demonstrates a reduced distribution gap between the source and target domains compared to these state-of-the-art methods, indicating its superiority in learning domain-invariant features. In contrast, the representation generated by iDAG exhibits a noticeable distinction between the source and target domains, indicating its weakness in tackling distribution shifts. For GMoE, there exist target features that are away from the source features, as indicated by the black circles. Besides, the representations produced by GMoE exhibit poor intra-class compactness. For VMamba, the ability to align features is inferior to our DGMamba, manifesting by the distant target features marked by the black circle, and the distinction between classes is not as clear as that in DGMamba.
Table 12: Comparison of state-of-the-art DG methods with our DGMamba with diverse backbones. The best results are bolded.
Method Backbone Params. Target domain Avg.($\uparrow$)
Art Cartoon Photo Sketch
iDAG [30] (ICCV’2023) ResNet50 25M 90.8 83.7 98.0 82.7 88.8
iDAG [30] (ICCV’2023) ResNet101 43M 89.0 84.9 98.3 84.7 89.2
GMoE [35] (ICLR’2023) DeiT-S 34M 89.4 83.9 99.1 74.5 86.7
GMoE [35] (ICLR’2023) DeiT-B 133M 91.0 84.0 99.3 82.7 89.2
DGMamba (ours) VMamba-T 31M 91.3 87.0 99.0 87.3 91.2
DGMamba (ours) VMamba-S 49M 94.1 87.8 99.6 89.0 92.6
DGMamba (ours) VMamba-B 88M 95.1 89.2 99.8 87.9 93.0
Performance with Diverse Backbones. To fully unleash the potential of our proposed DGMamba, we investigate the impact of utilizing larger backbones, i.e., stacking more Mamba layers or increasing the embedding dimension to facilitate capturing genuine features. Specifically, we conduct experiments on PACS utilizing VMamba-S and VMamba-B. VMamba-S comprises 4 blocks, each including 2, 2, 15, and 2 Mamba layers, with an embedding dimension of 96. VMamba-B maintains the same Mamba blocks and layers as VMamba-S, with an increased embedding dimension of 128. The generalization performances are concluded in Table 12, demonstrating that deeper Mamba architectures or larger embedding dimensions could enhance the model generalizability. Furthermore, our proposed DGMamba demonstrates superior generalization performance compared to CNN-based or ViT-based models, while maintaining comparable or fewer parameters. These results underscore the effectiveness of DGMamba in mitigating domain shifts.